Grade 12th passMechanics

A chain of mass M and length l is placed on a table with one sixth of it hanging freely from the table edge. What is the amount of work done to pull the chain on the table?

Orooba Asim
10 Years agoGrade 12th pass

2 Answers

Arohi
10 Years ago
First of all calculate linear mass density of chain that is equal to mass/length. next since 1/6 length is hanging the centre of mass will rise to a height of 1/12length on pulling the chain up.next work done =m*g*h=(1/6L*m/L)*g*1/12L=1/72 mgl.
Aryan Bharadwaj
4 Years ago
We know that:
W = mgl ÷ 2(n)²
 
Here substitute the value of n as 6 because it is given one-sixth in the question.
 
So, W = mgl ÷ 2(6)²
→ W = mgl ÷ (2 × 36)
→ W = mgl ÷ 72
 
Therefore the correct answer is \boxed{\bf{\dfrac{mgl}{72}}}